# WaveSP-Net: Learnable Wavelet-Domain Sparse Prompt Tuning for Speech Deepfake Detection

### v1

@@ 1 Introduction
Speech deepfake detection (SDD) is the task of identifying artificially generated or manipulated speech audio, distinguishing it from bonafide human speech. This capability is critical for protecting speaker verification systems from various attacks, including speech synthesis, voice conversion, and voice cloning. Remarkable progress in SDD has been made on both front-end features [1, 2, 3] and back-end models  [4, 5, 6], achieving promising detection results especially on intra-domain settings. However, generalization to diverse unseen domains remains a major challenge; real-world settings require generalization to new domains that may include unseen attacks, speech codecs, and audio compression formats [7].
As in other detection tasks, the choice of front-end features for SDD is critically important. Existing front-ends can be broadly categorized into digital signal processing (DSP) and self-supervised learning (SSL) based approaches, each offering distinct advantages for cross-domain generalization. On one hand, the former includes methods such as short-time Fourier transform, linear-frequency cepstral coefficients, and constant-Q transform  [8, 9, 10] aimed at capturing time-frequency characteristics using fixed transforms. On the other hand, modern data-driven SSL front-ends leverage foundational speech models such as XLS-R [11] and Wav2Vec 2.0 [12] to extract information-rich features. SSL front-ends are typically fine-tuned on the new domain [13, 14, 15, 16].
Fig. 1: Overview of the WaveSP-Net architecture. The figure illustrates five different XLSR-based front-end variants: (a) PT-XLSR, (b) FourierPT-XLSR, (c) WPT-XLSR, (d) WSPT-XLSR and Partial-WSPT-XLSR. The proposed WaveSP-Net (rightmost panel) integrates a Partial-WSPT-XLSR front-end (bottom right) with a Mamba-based classifier (top right).
While SSL front-ends typically achieve higher detection performance, they are computationally demanding and parameter-heavy, particularly with large SSL models with millions of parameters [17]. As a data-driven technique, they are also prone to overfitting. To address these challenges, parameter-efficient fine-tuning (PEFT) [18] has emerged as a practical solution. PEFT refers to a broad family of methods aimed at adapting a foundation model to new domains while keeping the number of parameters requiring updating small. For instance, [19] proposed intra-block and cross-block adapters to capture multi-level discriminative spoofing cues, whereas [20] integrated LoRA adapters into the self-attention heads of XLSR-AASIST [21], combined with meta-learning [22].
@@ 2.2 WSPT-XLSR & Partial-WSPT-XLSR
(1) Learnable Wavelet Decomposition (LWD). Learnable wavelet transforms have used earlier in other applications such as compression of neural networks [33], which is capable of dynamically adapting to different frequency domain signal characteristics. Inspired by this, we propose LWD. As illustrated in Fig. 1, from each of the layer-specific prompt token sets $\mathbf{P}_{k}$, we select the last $m$ tokens $\mathbf{P}_{k}^{(p-m+1:p)}$ and transform them into wavelet sparse prompt tokens $\mathbf{WSP}_{k}\in\mathbb{R}^{m\times d}$. In wavelet analysis [27], a signal is separated into two complementary components: a low-frequency part that captures the overall, coarse structure of the input, and a high-frequency part that captures the fine-grained detail information. This decomposition is performed using a pair of analysis filters, denoted by $F_{0}$ (low-pass) and $F_{1}$ (high-pass). While in DSP applications, these filters are typically selected from a set of preset ’library’ wavelets (e.g., Haar or Daubechies), in our model, they are learnable [33]; the filter coefficients are optimized during PT training, allowing the model to adaptively extract coarse and fine information that is most useful for the task.
@@ 2.3 WaveSP-Net
After fetching $\mathbf{WSP}_{k}$, we obtain the final prompt representation that integrates enhanced and untransformed tokens:
@@ 3.1 Dataset & Metrics
Our experiments use two benchmarks: Deepfake-Eval-2024 [28] and SpoofCeleb [29]. We train and evaluate each dataset separately following its official protocol. For Deepfake-Eval-2024 (DE24) [^2] https://huggingface.co/datasets/nuriachandra/Deepfake-Eval-2024, we follow [28] and preprocess the audio subset by chunking long clips into 4-second segments. For SpoofCeleb, we follow the established protocol [^3] https://www.jungjee.com/spoofceleb/. Performance is reported with EER, AUC, F1, and accuracy (ACC). We also report 95% parametric confidence intervals for EER following [36]: $\text{EER}\pm\sigma\cdot Z_{\alpha/2}$, where $Z_{\alpha/2}=1.96$, $\sigma=0.5\sqrt{\text{EER}(1-\text{EER})\left({n_{r}+n_{f}})/({n_{r}n_{f}}\right)}$, where $n_{r}$ and $n_{f}$ denote the number of real and fake samples, respectively.
@@ 3.2 Implementation Details
Table 2: Comparison with SOTA single systems on the Deepfake-Eval-2024 benchmark. The best results are in bold, and the second-best are underlined. The 95% parametric confidence intervals for EER are shown in parentheses. BCM denotes the Best Commercial Model. Params denotes trainable parameters.
@@
Table 3: Comparison with SOTA single systems on the SpoofCeleb benchmark. The best results are in bold. The 95% parametric confidence intervals for EER are shown in parentheses.
@@
4 Results and analysis
4.1 Framework with three novel XLSR variants Front-ends
@@ 4.3 Ablation & Parameter Sensitivity Experiments
Table 4 provides a detailed ablation study on the core components of the WaveSP-Net. Our results show that removing any core component leads to a notable performance degradation, with WDS causing the most significant drop by relatively 35.54% in EER. This highlights the critical role of the sparsity mechanism in filtering out noise. Additionally, replacing learnable wavelet filters with fixed ones also decreased performance, with a relative increase of 56.44% in EER, validating that learnable wavelet filters are effectively co-optimized with the back-end to learn discriminative features. We also perform hyperparameter sensitivity analysis on two key parameters: sparsity ratio and the number of wavelet sparse prompt tokens, as shown in Table 4. Experimental results indicate that the optimal WaveSP-Net configuration consists of learnable wavelet filters, a sparsity ratio of 0.1, and four wavelet sparse prompt tokens.
@@ 4.4 Visualization
Figure 2 presents a 2D t-SNE visualization of the DE24 test set. In Figs. 2(a) and (b), FourierPT-XLSR and WSPT-XLSR show significant overlap between real (blue) and fake (red) samples, indicating poor detection performance. In contrast, Partial-WSPT-XLSR, shown in Figure 2(c), displays distinct, tight, and highly isolated clusters with minimal overlap. The ideal separation demonstrates that WaveSP-Net effectively learns highly discriminative features by focusing on sparse, informative features in the wavelet domain.
@@ 5 Conclusion
This paper introduces WaveSP-Net, a novel audio deepfake detector that combines a Partial-WSPT-XLSR front-end with a bidirectional Mamba back-end. The core innovation lies in using learnable wavelet filters to enhance a sparse subset of prompt tokens adaptively. This approach proves to be a parameter-efficient and effective solution for SDD. Our experiments indicate that WaveSP-Net outperforms other SOTA single systems on two new and challenging benchmarks, Deepfake-Eval-2024 and SpoofCeleb, achieving SOTA performance with low trainable parameters. This successful integration of classical signal processing transforms into our architecture prompts us to reconsider their role. We believe these findings will inspire new approaches that combine hand-crafted acoustic features with the power of Large Language Models.

### v2

@@ 1 Introduction
Speech deepfake detection (SDD) is the task of identifying artificially generated or manipulated speech audio, distinguishing it from bonafide human speech. This capability is critical for protecting speaker verification systems from various attacks, including speech synthesis, voice conversion, and voice cloning. Remarkable progress in SDD has been made on both front-end features [37, 10, 38] and back-end models  [33, 35], achieving promising detection results especially on intra-domain settings. However, generalization to diverse unseen domains remains a major challenge; real-world settings require generalization to new domains that may include unseen attacks, speech codecs, and audio compression formats [20, 31].
As in other detection tasks, the choice of front-end features for SDD is critically important. Existing front-ends can be broadly categorized into digital signal processing (DSP) and self-supervised learning (SSL) based approaches, each offering distinct advantages for cross-domain generalization. On one hand, the former includes methods such as short-time Fourier transform, linear-frequency cepstral coefficients, and constant-Q transform  [24, 23, 25] aimed at capturing time-frequency characteristics using fixed transforms. On the other hand, modern data-driven SSL front-ends leverage foundational speech models such as XLSR [2] and Wav2Vec 2.0 [3] to extract information-rich features. SSL front-ends are typically fine-tuned on the new domain [16, 21, 32, 15, 34, 39, 17].
Fig. 1: Overview of the WaveSP-Net architecture. The figure illustrates five different XLSR-based front-end variants: (a) PT-XLSR, (b) FourierPT-XLSR, (c) WPT-XLSR, (d) WSPT-XLSR and Partial-WSPT-XLSR. The proposed WaveSP-Net (rightmost panel) integrates a Partial-WSPT-XLSR front-end (bottom right) with a Mamba-based classifier (top right). (FFT: Fast Fourier Transform; DWT: Discrete Wavelet Transform; IDWT: Inverse Discrete Wavelet Transform)
While SSL front-ends typically achieve better detection performance over conventional models, they are computationally demanding and parameter-heavy, particularly with large SSL models with millions of parameters [18]. As a data-driven technique, they are also prone to overfitting. To address these challenges, parameter-efficient fine-tuning (PEFT) [7] has emerged as a practical solution. PEFT refers to a broad family of methods aimed at adapting a foundation model to new domains while keeping the number of parameters requiring updating small. For instance, [29] proposed intra-block and cross-block adapters to capture multi-level discriminative spoofing cues, whereas [13] integrated LoRA adapters into the self-attention heads of XLSR-AASIST [11], combined with meta-learning [26].
@@ 2.2 WSPT-XLSR & Partial-WSPT-XLSR
(1) Learnable Wavelet Decomposition (LWD). Learnable wavelet transforms have used earlier in other applications such as compression of neural networks [28], which is capable of dynamically adapting to different frequency domain signal characteristics. Inspired by this, we propose LWD. As illustrated in Fig. 1, from each of the layer-specific prompt token sets $\mathbf{P}_{k}$, we select the last $m$ tokens $\mathbf{P}_{k}^{(p-m+1:p)}$ and transform them into wavelet sparse prompt tokens $\mathbf{WSP}_{k}\in\mathbb{R}^{m\times d}$. In wavelet analysis [19], a signal is separated into two complementary components: a low-frequency part that captures the overall, coarse structure of the input, and a high-frequency part that captures the fine-grained detail information. This decomposition is performed using a pair of analysis filters, denoted by $F_{0}$ (low-pass) and $F_{1}$ (high-pass). While in DSP applications, these filters are typically selected from a set of preset ’library’ wavelets (e.g., Haar or Daubechies), in our model, they are learnable [28]; the filter coefficients are optimized during PT training, allowing the model to adaptively extract coarse and fine information that is most useful for detecting deepfake speech.
@@ 2.3 WaveSP-Net
After computing $\mathbf{WSP}_{k}$, we obtain the final prompt representation that integrates both enhanced and untransformed tokens:
@@ 3.1 Dataset & Metrics
Our experiments use two benchmarks: Deepfake-Eval-2024 (DE24)  [6] and SpoofCeleb [12]. To evaluate deepfake detector generalization, we train and evaluate on each dataset separately following its official protocol. For Deepfake-Eval-2024  [^2] https://huggingface.co/datasets/nuriachandra/Deepfake-Eval-2024, we follow [6] and preprocess the audio subset by chunking long clips into 4-second segments. Spanning 88 web domains and 42 languages, DE24 includes audio samples with varying acoustic conditions, thereby subjecting the detector to strictly unseen attacks and complex distribution shifts. For SpoofCeleb, we follow the established protocol [^3] https://www.jungjee.com/spoofceleb/: the attacks included in the training are A01-A10, while the ones for evaluations are A15-A23. For more details, please refer to [12]. Performance is reported with EER, AUC, F1, and accuracy (ACC). We also report 95% parametric confidence intervals for EER following [4]: $\text{EER}\pm\sigma\cdot Z_{\alpha/2}$, where $Z_{\alpha/2}=1.96$, $\sigma=0.5\sqrt{\text{EER}(1-\text{EER})\left({n_{r}+n_{f}})/({n_{r}n_{f}}\right)}$, where $n_{r}$ and $n_{f}$ denote the number of real and fake samples, respectively.
@@ 3.2 Implementation Details
4 Results and analysis
4.1 Framework with Three Novel XLSR Variants Front-Ends
Table 2: Comparison with SOTA single systems on the Deepfake-Eval-2024 benchmark. The best results are in bold, and the second-best are underlined. The 95% parametric confidence intervals for EER are shown in parentheses. BCM denotes the Best Commercial Model.
@@ 4.1 Framework with Three Novel XLSR Variants Front-Ends
Table 3: Comparison with SOTA single systems on the SpoofCeleb benchmark. The best results are in bold. The 95% parametric confidence intervals for EER are shown in parentheses. Params denotes trainable parameters.
@@ 4.3 Ablation & Parameter Sensitivity Experiments
Table 4 provides a detailed ablation study on the core components of the WaveSP-Net. Our results indicate that removing any core component leads to a notable performance degradation, with WDS causing the most significant drop by relatively 35.54% in EER. This highlights the critical role of the sparsity mechanism in filtering out noise. Additionally, replacing learnable wavelet filters with fixed ones also decreased performance, with a relative increase of 56.44% in EER, validating that learnable wavelet filters are effectively co-optimized with the back-end to learn discriminative features. We also perform hyperparameter sensitivity analysis on two key parameters: sparsity ratio and the number of wavelet sparse prompt tokens, as shown in Table 4. Experimental results indicate that the optimal WaveSP-Net configuration consists of learnable wavelet filters, a sparsity ratio of 0.1, and four wavelet sparse prompt tokens.
@@ 4.4 Visualization
Fig. 2 presents a 2D t-SNE visualization of the DE24 test set. In Figs. 2(a) and (b), FourierPT-XLSR and WSPT-XLSR show significant overlap between real (blue) and fake (red) samples, echoing their detection performance. In contrast, Partial-WSPT-XLSR, shown in Fig. 2(c), displays distinct, tight, and highly isolated clusters with minimal overlap. This visualization demonstrates that WaveSP-Net effectively learns highly discriminative features by focusing on sparse, informative features in the wavelet domain.
@@ 5 Conclusion
This paper introduces WaveSP-Net, a novel speech deepfake detector that combines a Partial-WSPT-XLSR front-end with a bidirectional Mamba back-end. The core innovation lies in using learnable wavelet filters to enhance a sparse subset of prompt tokens adaptively. This approach turns out to be a parameter-efficient and effective solution for SDD. Our experiments indicate that WaveSP-Net outperforms other SOTA single systems on two new and challenging benchmarks, Deepfake-Eval-2024 and SpoofCeleb, achieving SOTA performance with low trainable parameters. This successful integration of classical signal processing transforms into our architecture prompts us to reconsider their role. We believe these findings will inspire new approaches that combine hand-crafted acoustic features with the power of large language models.
6 Acknowledgment
This work was supported by the Finnish Doctoral Program Network in Artificial Intelligence (AI-DOC), project “Explainable Speech Deepfake Characterization” (Decision No. VN/3137/2024-OKM-6), and partially by the Research Council of Finland, project “SPEECHFAKES” (Decision No. 349605).